OpenStudy (anonymous):

3x - 3y = 12 4y - 8x = -16

DivineSolar (divinesolar):

For the first one 3x - 3y what do you think your supposed to do first?

DivineSolar (divinesolar):

I have to walk you through this.

OpenStudy (anonymous):

Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.

DivineSolar (divinesolar):

Ok so the first step is add 3y to both sides.

DivineSolar (divinesolar):

Now what do you get when you do that?

DivineSolar (divinesolar):

Hello?

OpenStudy (anonymous):

3x-3y=12 3x=12+3y x=(12+3y)/3 substitute this in the equation 4y-8x=-16 4y-8(12+3y)/3=-16 4y-(96/3)-(24y/3)=-16 4y-32-8y=-16 -4y=-16+32 -4y=16 y=-16/4 y=-4 x=(12-12)/3=0

DivineSolar (divinesolar):